OpenStudy (anonymous):

PLEASE HELP! :( The curve C has equation y={a(x-a)^2}/x^2-4a^ where a is a positive constant. Show that C has one maximum point and one minimum point and find their coordinates.

myininaya (myininaya):

Find y' Set y' equal to 0 Solve y'=0 for x to get the critical numbers. Then we will determine if the critical numbers are max or mins or neither.

OpenStudy (anonymous):

\[y={a(x-a)^2}\div x^2-4a^2\]

OpenStudy (anonymous):

I couldnot solve y'=0. I TRIED! :(

myininaya (myininaya):

What did you get for y'

myininaya (myininaya):

Oh and we also need to find where y' does not exist

myininaya (myininaya):

Use quotient rule.

OpenStudy (anonymous):

a(x-a)^2(2x-8a)-(x^2-4a^2)(2ax-2a^2)=0

myininaya (myininaya):

hmmm.... You know a is a constant right?

OpenStudy (anonymous):

Yep!

myininaya (myininaya):

Derivative of a is 0. Derivative of any constant *a is 0.

myininaya (myininaya):

a^2 is also a constant since a is a constant (a^2)'=0

OpenStudy (anonymous):

Oops. Now I get it. O_o

myininaya (myininaya):

And don't forget the quotient rules is : \[y'=\frac{ (\text{ derivative of top }) \cdot (\text{ bottom } ) - ( \text{ derivative of bottom } ) \cdot ( \text{ top } )}{(bottom)^2}\]

OpenStudy (anonymous):

Can you please show me what exactly y' would be? All these 'a' and brackets have confused me. :(

OpenStudy (anonymous):

@experimentX : Can you please help me with this question?

myininaya (myininaya):

\[y=\frac{a(x-a)^2}{x^2-4a^2}\] \[ \text{ What is } [a(x-a)^2]' \text{ equal to? }\] \[ \text{ What is } [x^2-4a^2]' \text{ equal to?}\]

myininaya (myininaya):

For my first question, a is a constant multiple. We can just bring that down and look at differentiating (x-a)^2 by using chain rule.

myininaya (myininaya):

For my second question, 4a^2 is just a constant so we only need to look at differentiating x^2 by use of the power rule.

myininaya (myininaya):

When we get done, we will plug in the bottom, the top, the derivative of the bottom, and the derivative of the top into the following formula I gave you earlier: \[y'=\frac{ (\text{ derivative of top }) \cdot (\text{ bottom } ) - ( \text{ derivative of bottom } ) \cdot ( \text{ top } )}{(bottom)^2} \]

myininaya (myininaya):

So so far we have: \[y'=\frac{ (\text{ derivative of top }) \cdot (x^2-4a^2) - ( \text{ derivative of bottom } ) \cdot ( a(x-a)^2 )}{(x^2-4a^2)^2} \]

myininaya (myininaya):

I'm just asking you now to give me the derivative of \[ (a(x-a))^2 \text{ and } (x^2-4a^2) \].

OpenStudy (anonymous):

For (a(x-a))^2, it would be 2x-2a. For (x^2-4a^2), it would be 2x.

myininaya (myininaya):

For the first one, couldn't you just leave it as 2(x-a) but don't forget to bring down that a in front. :) So you would actually have what for the first one?

OpenStudy (anonymous):

Would it be 2a(x-a)? :O

myininaya (myininaya):

Yep so we have \[y'=\frac{ (\text{ derivative of top }) \cdot (x^2-4a^2) - ( \text{ derivative of bottom } ) \cdot ( a(x-a)^2 )}{(x^2-4a^2)^2} \] = \[y'=\frac{ ( 2a(x-a) ) \cdot (x^2-4a^2) - ( 2x ) \cdot ( a(x-a)^2 )}{(x^2-4a^2)^2}\] = \[=\frac{2a(x-a)(x^2-4a^2)-2xa(x-a)^2}{(x^2-4a^2)^2}\]

myininaya (myininaya):

Now what factors in the numerator do the two terms on top have in common?

OpenStudy (anonymous):

2a and x-a.

myininaya (myininaya):

Yes. :) So factor that out.

myininaya (myininaya):

\[2a(x-a)(x^2-4a^2)-2xa(x-a)^2\] = \[2a(x-a)[ ? - ?]\]

OpenStudy (anonymous):

2a(x-a){(x^2-4a^2)-x(x-a)}

myininaya (myininaya):

Yep You will have \[\frac{2a(x-a)[(x^2-4a^2)-x(x-a)]}{(x^2-4a^2)^2}\] You can simplify what you have in brackets above though.

myininaya (myininaya):

\[\frac{2a(x-a)(x^2-4a^2-x^2+xa)}{(x^2-4a^2)^2}\] You have like terms. :)

myininaya (myininaya):

Do you see the like terms to combine?

OpenStudy (anonymous):

Yep. I'll end up with 2a(x-a)(xa-4a^2). Right?

myininaya (myininaya):

Ok great now set equal factor equal to zero You are almost there :)

myininaya (myininaya):

a is a positive constant so 2a can never be zero so set x-a equal to zero and set xa-4a^2=0 Solve both for x to find the critical numbers :)

OpenStudy (anonymous):

So I'll have x=a and x=4a? What do I do now?

myininaya (myininaya):

It said find the coordinates so find what y is if x=a and find what y is if x=4a

OpenStudy (anonymous):

If x=a, then y=0. If x=4a, y=9a^3. Is it correct?

OpenStudy (anonymous):

No. Sorry. For x=4a, y is undefined.

myininaya (myininaya):

Your first one is right. Your second one is wrong.

OpenStudy (anonymous):

So y is not undefined?

OpenStudy (anonymous):

Oops. I found out my mistake. For x=4a, y=3a/4. Thank you soooooo muchhhhh @myininaya ! :D

myininaya (myininaya):

Great. :)

OpenStudy (anonymous):

One more thing, how do I show which is the minimum point and which is maximum? Is it by the second derivative?